Find the equation of the line which passes through the point (1, √ 3) and makes an angle of 30◦ with the line x− √ 3y + √ 3 = 0


1
Expert's answer
2021-06-07T13:29:15-0400

Write the equation of the line in slope-intercept form


"y=mx+b"

Line 

"x-\sqrt{3}y+\sqrt{3}=0"

"y=\dfrac{1}{\sqrt{3}}x+1"

"slope_1=m_1=\dfrac{1}{\sqrt{3}}"

The nagle between two lines


"\tan \theta=\pm\dfrac{m_2-m_1}{1+m_1m_2}"

Given "\theta=30\degree"


"\tan 30\degree=\pm\dfrac{m_2-m_1}{1+m_1m_2}"

"\dfrac{1}{\sqrt{3}}=\pm\dfrac{m_2-\dfrac{1}{\sqrt{3}}}{1+\dfrac{1}{\sqrt{3}}m_2}"

"\sqrt{3}+m_2=\pm(3m_2-\sqrt{3})"

Let


"\sqrt{3}+m_2=-(3m_2-\sqrt{3})"

"m_2=0"

The equation of the line which passes through the point "(1, \sqrt{3})" is


"y=\sqrt{3}"


Let


"\sqrt{3}+m_2=(3m_2-\sqrt{3})"


"m_2=\sqrt{3}"

Point "(1, \sqrt{3})"

"\sqrt{3}=\sqrt{3}(1)+b=>b=0"

The equation of the line which passes through the point "(1, \sqrt{3})" is


"y=\sqrt{3}x"

"y=\sqrt{3}" or "y=\sqrt{3}x"